OpenStudy (anonymous):

Find the inverse function of y = 2 log(3x) + log(5)

OpenStudy (anonymous):

log 5*9x^2 = y 45x^2 = 10^y x^2 = 10^y / 45 x = [10^(y/2)] / sqrt45 inverse function f(-1) = [ 10^(x/2)] / sqrt45

OpenStudy (anonymous):

thank you jimmy!

OpenStudy (anonymous):

yw